OpenStudy (anonymous):

how to find instantaneous velocity with the function s(t)=4t+1 at t=2 seconds

OpenStudy (anonymous):

its velocity is not changing and it is always 4m/s

OpenStudy (anonymous):

So, I believe instantaneous velocity is the first derivative of the function. \[s \prime (t) = 4\] Meaning that there is no place to plug in t=2. So instantaneous velocity is static and should be 4.

OpenStudy (anonymous):

where did you get the 4?

OpenStudy (anonymous):

differenciate s(t) wrt time

OpenStudy (anonymous):

You need to find the first derivative of of s(t)= 4t+1. The derivative of 4t is 4(1) and the derivative of 1 is 0. s'(t)= 4(1)+0 = 4

OpenStudy (anonymous):

But doesn't the problem strep that t=2?

OpenStudy (anonymous):

Also if the function is 4t+1 shouldn't plugging 1 in be 5?

OpenStudy (anonymous):

No, because when you are looking for instantaneous velocity you need to plug t into the derivative of the original equation. However, in this case in the derivative equation there is no longer a t.

OpenStudy (anonymous):

Basically, you were given s(t)=4t+1 at t=2. The next step is to find the derivative of s(t)= 4t+1. s'(t) = 4 Now you plug t=2 into s'(t) = 4. However, there is no longer a t to plug into so the answer is just 4.

OpenStudy (anonymous):

Ok, so how did you find the derivative? did you use the slope equation?

OpenStudy (anonymous):

No, I did it in a way you might not have learned yet. But let me know if I'm wrong. Using the difference quotient: \[\lim_{h \rightarrow 0} \frac{ (4(t-h)+1)-(4t+1) }{ h }\] can you solve for the derivative from there?

OpenStudy (anonymous):

The way I did it was knowing these rules: d/dx (f(x)+g(x)) = f'(x)+g'(x) (derivative of sums is sum of derivative) d/dx cf(x) = cf'(x) (when taking the derivative of a constant, take out the constant and put it back after finding the derivative of f(x). d/dx x = 1 (derivative of a variable by itself is 1) d/dc c = 0 (derivative of a constant is 0)

OpenStudy (anonymous):

In the last line the equation should be d/dx c= 0. Sorry for the typo. Hope I helped clarify a bit :)

OpenStudy (anonymous):

Yeah, I didn't learn that, thanks anyway!